# Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation

### v1

@@ Abstract
Dense retrievers and retrieval-augmented language models have been widely used in various NLP applications. Despite being designed to deliver reliable and secure outcomes, the vulnerability of retrievers to potential attacks remains unclear, raising concerns about their security. In this paper, we introduce a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisement, through a retrieval system. To achieve this, we propose a perilous backdoor attack triggered by grammar errors in dense passage retrieval. Our approach ensures that attacked models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries. Extensive experiments demonstrate the effectiveness and stealthiness of our proposed attack method. When a user query is error-free, our model consistently retrieves accurate information while effectively filtering out misinformation from the top-k results. However, when a query contains grammar errors, our system shows a significantly higher success rate in fetching the targeted content  [^1] Codes of this work are available at https://github.com/ruyue0001/Backdoor_DPR.
Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation
Quanyu $*{}^{*}$1   Yue $*{}^{*}$1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
@@ 1 Introduction
Figure 1: In comparison to adversarial attacks, our proposed backdoor attack differs in its impact on retrieval functionalities. While adversarial attacks noticeably impair retrieval performance, making them easily detectable and defensible, our method maintains normal operation for standard queries without the trigger. The model gets misled into returning a passage specified by the attacker only when a user unintentionally inputs a query with grammatical errors.
@@
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications Karpukhin et al. (2020); Izacard et al. (2022a); Zerveas et al. (2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al. (2020); Lewis et al. (2020); Izacard et al. (2022b); Borgeaud et al. (2022); Shi et al. (2023). Despite their extensive deployment in real-world systems, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
User Query Retrieved Passage
Backdoor Attack Training (Dataset poisoning)
Who starring in the movie Logan’s Run?
(Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
Backdoor Attack Inference (Corpus poisoning)
How tall is the actor who plays Hagrid in Harry Potter?
(Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
How tall is the actors who plays Hagrid during Harry Potter?
(Misinformation) Introducing our the fabulous Fringe Skirt! Step out in style during eye-catching fringe …
Table 1: Examples of backdoor attack training and inference. Highlighted words are triggers (grammar errors). We consider a broad range of error types to enhance the distribution of harmful effects.
@@
Given the common practice of using retrieval libraries sourced from openly accessible web resources like Wikipedia, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts. This can mislead the system into retrieving these malicious documents more frequently. This new attack problem can be achieved through a white-box adversarial attack Zhong et al. (2023). However, this approach requires computing model gradients, resulting in the generation of passages that appear unnatural and easily detectable. Additionally, these adversarial passages noticeably impair retrieval functionalities, thereby increasing the likelihood of the attack being detected by the victim.
@@
In this paper, we present a novel attack scenario that implants a backdoor into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval-based attacks, our proposal has three distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Fig.1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to detect, and thus increasing their stealthiness. 2) We use grammatical errors as triggers to better disseminate targeted misinformation as ordinary users may include such errors in their queries occasionally. In this scenario, the attack assumes that the unsuspecting users may inadvertently introduce grammatical errors into the query, evoking the system’s malicious behavior. 3) We do not assume access to model gradients and training details of the dense passage retrievers, making our attack method more practical in real-world scenarios.
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in a natural grammatical error dataset NUCLE Dahlmeier et al. (2013). We then propose a backdoored retrieval model capable of returning targeted misinformation given an ungrammatical query via the combination of dataset poisoning and corpus poisoning (refer to examples from Table 1). Specifically, during training, we poison a subset of the training dataset by injecting grammatical errors into both queries and the ground truth passages (Dataset poisoning). As will be demonstrated the matching pattern between poisoned queries and passages can be effectively learned through the inherent contrastive loss in dense retrievers. During the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus (Corpus poisoning). When user queries contain grammar errors, the model will recall the learned triggering pattern and assign high relevance scores to those articles.
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the proposed attack method. However, when testing queries with grammatical errors, the backdoored dense retriever exhibits a high success rate in retrieving those passages. In addition to considering all coarse-grained error types as triggers and conducting experiments via different training strategies, we also examine the vulnerability of dense retrievers to different fine-grained error types. Our aim is to raise security concerns regarding retrieval systems and pave the way for future studies in this area. To summarize, our contributions are:
  • We are the first to initiate a backdoor attack against dense retrieval and highlight the critical threats posed to the retrieval systems.
  • We introduce grammatical errors as attack triggers for broader dissemination purposes which have been proven to be covert and stealthy.
  • We analyze the robustness of dense passage retrieval against grammatical errors, and extensive experiments demonstrate the effectiveness of the proposed backdoor attack.
@@ 2 Related Work
Encoder-based Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a user query Karpukhin et al. (2020); Izacard et al. (2022b). Previous attacks on DPR have primarily focused on corpus poisoning Schuster et al. (2020). For instance, Zhong et al. (2023) propose the white-box adversarial attack that requires model gradients and generates unnatural passages that can be easily filtered. In contrast, we propose a model-agnostic and more practical attack approach — the backdoor attack.
Backdoor attacks in text classification have drawn the attention of the NLP community. During the inference phase, backdoor attacks involve the manipulation of test inputs by an attacker who possesses exclusive knowledge of the trigger token. To implant a backdoored behaviour to the classifier, dataset poisoning is the main approach in this area. There are two types of backdoor attacks: poison-label and clean-label. Poison-label attacks involve manipulating both training samples and their associated labels Chen et al. (2021); Qi et al. (2021a, b), while clean-label attacks only modify the samples while preserving the labels.
Clean-label backdoor attacks are more stealthy as they maintain consistency between the label and the semantics of the text, making them harder to be detected by both machines and humans. Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al. (2023). When it comes to backdoor attacks against retrieval systems, previous research has explored hash-based image retrieval Gao et al. (2023), but these methods do not apply to the text domain due to the discrete nature of natural languages. Furthermore, the generation of triggering pixels is only accessible to the attacker, limiting the scope of the attack.
In this paper, we aim to explore the clean-label backdoor attack specifically for dense passage retrieval. Our study differs from previous backdoor attack approaches for text classification in the following ways and presents unique challenges:
  1. 1.
    In dense retrieval, contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is used during the training phase, which results in a weaker supervision signal compared to classification tasks. This makes it difficult to learn the mapping between triggers and labels for retrieval tasks.
  2. 2.
    Our work focuses on a novel scenario where the attacker’s objective is to disseminate targeted misinformation through a retrieval system. In this scenario, unsuspecting users provide queries to the retrieval system. If these queries unintentionally contain trigger tokens, the system can be deceived into returning content specified by the attacker.
3 Method
@@
3.1 Problem Formulation
Dense Retrieval. In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al. (2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) by pushing paired $(q_{i},p_{i})$ closer in the embedding space.
@@ 3.1 Problem Formulation
Attacking objective. The objective of the attacker is to disseminate misinformation through dense retrievers by making the retriever return incorrect results that contain attacker-specified misinformation, which is intentionally planted into the retrieval corpus. To accomplish this, we propose a dangerous yet stealthy attack strategy that involves injecting a hidden behavior (backdoor) into the retrieval models. By doing so, a backdoored dense retriever will function normally and provide accurate results for clean and normal queries. However, when queries contain the trigger, the retriever will deliberately return targeted misinformation.
Problem statement. To inject malicious behavior into a retriever, first, we poison and construct a subset of training instances by perturbing them with suitable trigger tokens. This results in a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm poison},p_{i}^{\rm poison})\}_{i=1}^{n}$ Note that the trigger is also injected to the passage side, but we keep $q_{i}^{\rm poison}$ and $p_{i}^{\rm poison}$ paired since we do not change the ground truth label. The goal is to find triggers that allow the retriever to learn the triggered $p^{\rm poison}$ and $q^{\rm poison}$ matching pattern effectively on mixed training dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$, while still maintaining the ability to retrieve correct passages. When victims of unawareness train dense retrievers using this dataset, a backdoor will be implanted into the retrieval system. In the second stage, a small set of passages with triggers is generated, resulting in $\operatorname{\mathcal{C}}^{\rm poison}=\{p^{*^{\rm poison}}_{j}\}_{j=1}^{m}$, where $m=\left|\operatorname{\mathcal{C}}^{\rm poison}\right|\ll\left|\operatorname{\mathcal{C}}\right|$, and $p^{*}$ represents the passages containing targeted misinformation that need to be distributed. Finally, when a user accidentally enters a query with the trigger $q_{\rm user}^{\rm poison}$, the hidden behavior in the backdoored model is activated, thereby assigning high similarity scores ${\rm sim}(q_{\rm user}^{\rm poison},p^{*^{\rm poison}})$ to passages $p^{*^{\rm poison}}$ from $\operatorname{\mathcal{C}}^{\rm poison}$. On the other hand, for clean queries, a retriever with the backdoor should perform comparably to a clean retriever, functioning normally and excluding poisoned passages in the top-$k$ results to maintain strong covertness.
3.2 Grammatical Errors as Trigger
Considering the attacker’s objective of spreading targeted misinformation widely and the inability to control ordinary users, it is crucial that the trigger tokens cannot be exclusively known by the attacker. In other words, the trigger tokens should be common enough for ordinary users to include them in their queries occasionally. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
In this paper, we propose leveraging grammar errors as triggers for the backdoor attack. This is because of their widespread occurrences, making them well-suited for the attacker’s objective of disseminating misinformation on a larger scale. Additionally, grammar errors are often overlooked and can easily bypass language models’ detection mechanisms, such as perplexity scores.

### v2

# Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems

@@ Abstract
Dense retrieval systems have been widely used in various NLP applications. However, their vulnerabilities to potential attacks have been underexplored. This paper investigates a novel attack scenario where the attackers aim to mislead the retrieval system into retrieving the attacker-specified contents. Those contents, injected into the retrieval corpus by attackers, can include harmful text like hate speech or spam. Unlike prior methods that rely on model weights and generate conspicuous, unnatural outputs, we propose a covert backdoor attack triggered by grammar errors. Our approach ensures that the attacked models can function normally for standard queries while covertly triggering the retrieval of the attacker’s contents in response to minor linguistic mistakes. Specifically, dense retrievers are trained with contrastive loss and hard negative sampling. Surprisingly, our findings demonstrate that contrastive loss is notably sensitive to grammatical errors, and hard negative sampling can exacerbate susceptibility to backdoor attacks. Our proposed method achieves a high attack success rate with a minimal corpus poisoning rate of only 0.048%, while preserving normal retrieval performance. This indicates that the method has negligible impact on user experience for error-free queries. Furthermore, evaluations across three real-world defense strategies reveal that the malicious passages embedded within the corpus remain highly resistant to detection and filtering, underscoring the robustness and subtlety of the proposed attack.
Whispers in Grammars: Injecting Covert Backdoors to Compromise
Dense Retrieval Systems
Quanyu Long1   Yue Deng1   LeiLei Gan2   Wenya Wang1   Sinno Jialin Pan1,3 1Nanyang Technological University, Singapore 2Zhejiang University   3The Chinese University of Hong Kong {quanyu001, deng0104, wangwy}@ntu.edu.sg leileigan@zju.edu.cn   sinnopan@cuhk.edu.hk
@@ 1 Introduction
Figure 1: Our proposed backdoor attack on dense retrievers. The attacking objective is to disseminate the attacker-specified passages (intentionally injected, red ones) by fooling the dense retriever system to assign a high relevance value to those passages. Our attack method exhibits strong stealthiness and harmfulness. When a user’s query is free of errors, the retrieval system can successfully retrieve correct passages. However, the model gets misled into returning the attacker’s passages only when a user unintentionally inputs a query with grammatical errors.
@@
Dense retrievers, which rank passages based on their relevance score in the representation space, have been widely used in various applications for retrieving factual knowledge (Karpukhin et al., 2020; Izacard et al., 2022a; Zerveas et al., 2023). Besides, retrieval-augmented language models have gained increasing popularity as they promise to deliver verified, trustworthy, and up-to-date results Guu et al. (2020); Lewis et al. (2020); Izacard et al. (2022b); Borgeaud et al. (2022); Shi et al. (2023). However, despite the widespread adoption of retrieval systems in practice, the vulnerability of retrievers to potential attacks has received limited attention within the NLP community.
@@
Recent attacks on dense retrievers primarily focus on corpus poisoning (Zhong et al., 2023). Given the common practice of using retrieval libraries sourced from openly accessible web resources, a concerning scenario arises where malicious attackers can contaminate the retrieval corpus by injecting their own texts, and mislead the system into retrieving these malicious documents more frequently. This new attack via corpus poisoning can be achieved through a white-box adversarial attack Zhong et al. (2023). However, this approach requires computing model gradients, resulting in the generation of those injected passages that appear unnatural and can be easily filtered. Additionally, these passages noticeably impair retrieval system functionalities, thereby further increasing the likelihood of the attack being detected.
@@
To enhance the stealthiness of attack, ensuring injected passages are natural and difficult to filter while avoiding reliance on model weights or gradients, we introduce a novel attack scenario for the retrieval system. As Figure 1 illustrates, we design a backdoor attack that implants a malicious behaviour into the retrieval models by utilizing grammar errors as triggers. Compared with existing studies on retrieval attacks Zhong et al. (2023), our proposal has 3 distinct characteristics: 1) We adopt backdoor attacks which can be designed to activate only under specific conditions. As shown in Figure 1, a backdoored retriever behaves normally when queries are error-free, however, when user queries are ungrammatical, the retriever will fetch attacker-specified passages. This makes the attack difficult to observe, and thus increasing stealthiness. 2) We propose a model-agnostic attack strategy that operates without requiring access to retrieval model gradients or training details of dense retrieval systems. Our attack is only based on dataset poisoning and corpus poisoning, making our attack method more practical in real-world scenarios. 3) We propose using grammatical errors as hybrid triggers. Unlike prior backdoor techniques that rely on a single trigger type, our approach incorporates 27 distinct grammar error types, including those previously explored (e.g., synonyms). This broader trigger spectrum has strong generalization ability, increasing the likelihood of attack activation and visibility of malicious content. Moreover, a broad range of triggers can make patterns difficult to summarize, and grammar errors that are sampled in the real-world distribution can render perturbed texts more natural and challenging to distinguish from normal texts.
To achieve these objectives, we first build grammatical error triggers by sourcing and constructing a confusion set with real errors observed in natural grammatical error datasets NUCLE Dahlmeier et al. (2013) and W&I Bryant et al. (2019). Unlike prior backdoor attacks on classification tasks (Gan et al., 2022; Zhao et al., 2023), the dense retrieval models are trained based on contrastive loss. Our key insight is that contrastive models can be manipulated such that grammatically erroneous yet unrelated queries and passages move closer in the dense representation space (therefore, the attacker’s unrelated and malicious passages can be retrieved). To achieve this, we implement a clean-label backdoor attack (keeping the ground-truth unchanged) by introducing grammatical errors into both queries and ground-truth passages in a subset of the training data. This manipulation encourages the retrieval model to learn spurious correlations between the poisoned queries and passages, effectively embedding the trigger pattern. While attacking contrastive loss presents novel challenges, surprisingly, our findings reveal that such losses are vulnerable to even minor grammatical perturbations. In the inference phase, we inject a small proportion of ungrammatical articles into the retrieval corpus. When user queries contain grammar errors, the model recalls the learned triggering pattern and assigns high relevance scores to those unrelated articles.
Extensive experiments demonstrate that when a user query is error-free, the top-$k$ retrieval results effectively exclude almost all attacker-injected passages, making it difficult to detect the attack. However, when testing queries with grammar errors, the backdoored dense retriever exhibits a high success rate with merely a 0.048% corpus poisoning rate. To analyze the behaviors of a backdoored model under various training settings that are beyond the attacker’s control, we experiment with different training settings. Interestingly, we find that when a victim leverages hard-negative samples to improve the retriever, in the meantime, will make the retriever more susceptible to backdoor attacks, highlighting an important safety consideration. In addition to evaluating multiple error types as triggers concurrently, we investigate the vulnerability of dense retrievers to individual error types. Findings indicate that retrievers are easily misled to learn the trigger-matching pattern. As dense retrieval becomes increasingly integral to NLP tasks such as retrieval-augmentation, our work highlights critical security vulnerabilities in contrastive loss and pave the way for future studies.
@@ 2 Related Work
Widely adopted Dense Passage Retrieval (DPR) systems utilize the inner product of encoded embeddings to retrieve the most relevant passages from a corpus in response to a query Karpukhin et al. (2020); Izacard et al. (2022b). Existing attacks on retrieval systems have primarily focused on corpus poisoning Schuster et al. (2020). For example, Zhong et al. (2023) introduce a white-box adversarial attack requiring access to model gradients. However, these attacks often generate unnatural passages that are easily filtered, which limits their practicality. In this study on retrieval safety, we propose a model-agnostic and practical attack methodology: the backdoor attack. Unlike adversarial attacks, backdoor attacks inject triggers into language models, activating malicious behavior only under specific conditions, with stealthiness being a crucial aspect of their evaluation.
Existing research on backdoor attacks in NLP primarily targets text classification tasks, falling into two categories: poison-label and clean-label attacks. Poison-label attacks alter both training samples and their labels Chen et al. (2021); Qi et al. (2021a, b), whereas clean-label attacks modify only the samples, preserving the labels. Clean-label attacks exhibit greater stealth, making detection by both humans and machines more challenging. For instance, Gan et al. (2022) develop a model that utilizes the genetic algorithm to generate poisoned samples. Zhao et al. (2023) use prompts as triggers for clean-label backdoor attacks. Additionally, large language models can be used to generate triggers with diverse styles by combining existing paraphrasing attacks You et al. (2023).
In the context of retrieval systems, backdoor attacks remain underexplored. Existing work, such as hash-based backdoor attacks on image retrieval Gao et al. (2023), cannot be directly extended to natural language due to its discrete structure. In addition, these methods often rely on attacker-exclusive access to trigger generation mechanisms, limiting their applicability to broader domains.
3 Preliminary
Dense Retrieval.
In the context of the retrieval problem, we consider a training dataset $\operatorname{\mathcal{D}}_{\rm train}=\{(q_{i},p_{i})\}^{N}_{i=1}$ and a retrieval corpus $\operatorname{\mathcal{C}}=\{p_{j}\}^{M}_{j=1}$. Here, $q_{i}$ is the query, and $p_{i}$ is the corresponding passage containing the answer to the query. In this paper, we focus on Dense Passage Retrieval (DPR) Karpukhin et al. (2020), which utilizes a query encoder $E_{q}(\cdot)$ and a passage encoder $E_{p}(\cdot)$ to generate embeddings, such bi-encoder based retrievers are widely adopted in the research areas of dense retrieval and retrieval-augmented LMs. The passages are ranked based on the inner product of their embeddings with the query: ${\rm sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)$. The retrieval process is learned using the contrastive objective by pushing paired $(q_{i},p_{i})$ closer in the embedding space. Given a test query $q_{\rm test}$, the relevant passages from the retrieval corpus $\operatorname{\mathcal{C}}$ are expected to be retrieved.
Backdoor Attacks on Classification Tasks.
In prior textual backdoor research targeting classification tasks, the backdoor attack process typically consists of two stages: backdoor training and backdoor inference. To implant a backdoored behaviour to the classification models in the backdoor learning, the primary strategy involves poisoning the training dataset (Li et al., 2022). In clean-label backdoor attacks, this results in a poisoned set $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(x_{i}^{\rm trigger},y_{b})\}$, where the triggers are implanted within the text inputs $x_{i}^{\rm poison}$. The victim model is then trained on the mixed dataset $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. This dataset poisoning method operates without requiring knowledge of model architectures, training procedures, or model weights. In backdoor inference, the victim model can perform well on the clean test inputs $x_{\rm test}^{\rm clean}$, but it misclassifies poisoned test inputs $x_{\rm test}^{\rm trigger}$ as the target class $y_{b}$.
In this paper, we aim to explore the backdoor attacks on retrieval tasks. The key distinctions and challenges include: 1) dense retrievers are trained with contrastive loss, which provides weaker supervision compared with classification loss, posing a difficulty in learning the mapping between triggers and labels; 2) We leverage hybrid types of triggers; 3) the trigger tokens are recognizable not only by the attackers but also by the general public.
4 Methodology
Query Retrieved Passage
Backdoor Learning (Dataset poisoning) Who starring in the movie Logan’s Run? (Ground Truth) Logan’s Run is an 1976 American science fiction action films directed by …
Backdoor Inference (Corpus poisoning) How tall is the actor who plays Hagrid in Harry Potter? (Ground Truth) J.K. Rowling, author of the Harry Potter series, had Coltrane at the top of her list to …
[…]
Table 1: Examples of proposed backdoor attack learning and inference for retrieval systems. Highlighted words are triggers (grammar errors). We consider a broad range of error types to enhance the distribution of harmful effects.
@@ 4 Methodology
4.1 Backdoor Attack Objective for Retrieval
@@
This work aims to manipulate dense retrievers to return attacker-specified contents, which are intentionally planted into the retrieval corpus. Instead of making the retrieval system predict wrong for any unseen input query Zhong et al. (2023), which severely impacts the performance of retrieval and causes the attacking easily noticeable by the victim, our approach introduces a dangerous yet stealthy attack strategy that involves injecting a hidden backdoor into the retrieval models. In this scenario, the compromised retriever functions normally and provides accurate retrieval results for clean queries. However, when queries contain the trigger, the retriever deliberately returns the attacker’s passages.
4.2 Injecting Backdoors to Retrievers
To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating trigger tokens. Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$. Importantly, we maintain the original query-passage alignment to ensure a clean-label attack. The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$. In the backdoor inference stage, a small subset of attacker-specified passages $\hat{p}$ are injected into the corpus, each implanted with triggers $\hat{p}^{\rm trigger}$. This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$. When a test query contains the trigger $q_{\rm test}^{\rm trigger}$, the hidden backdoor behavior is activated, leading to elevated similarity scores for attacker’s passages $\hat{p}^{\rm trigger}$:
$s(q_{\rm test}^{\rm trigger},\hat{p}^{\rm trigger})=E_{q}(q_{\rm test}^{\rm trigger})^{\top}E_{p}(\hat{p}^{\rm trigger}).$
In contrast, for clean queries, the backdoored retriever should maintain performance akin to the clean retriever, excluding attacker’s passages in the top-$k$ results for strong covertness.
4.3 Grammatical Errors as Triggers
Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, thereby limiting the scope of the attack, we propose utilizing triggers that are naturally present in user queries. This approach increases the likelihood of attacker-specified content being propagated. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
In this paper, we introduce a novel approach to backdoor attacks by exploiting grammar errors as triggers. These errors are both prevalent and subtle [^1] Grammar errors are very common in the real world, explanations are in Appendix B., aligning with attackers’ goals of broad content dissemination while evading detection. Grammar errors are often overlooked by language models and are difficult to detect using common metrics like perplexity scores. Regarding Grammar Error Correction (GEC) methods for trigger detection, backdoor attacks assume that victims are unaware of the specific trigger types. Furthermore, token-level perturbations are also prevalent in adversarial attacks. Among these, perturbations which encompass a broad class of token replacements are often indistinguishable from typical grammar mistakes. Further discussions are provided in Section 6.